OpenStudy (anonymous):

Find the indefinite integral. (Remember to use ln |u| where appropriate.) (x^3 − 3x^2 + 5x − 12)/(x^2 + 4) dx

OpenStudy (anonymous):

its an indefinite integral but im lost on how to even solve it

myininaya (myininaya):

have you divided?

myininaya (myininaya):

x-3 ----------------------------- x^2+4 | x^3 -3x^2 +5x -12 -(x^3 +4x) ----------------------- -3x^2+x-12 -(-3x^2 -12) -------------- x

myininaya (myininaya):

so i got \[\frac{x^3-3x^2+5x-12}{x^2+4}=x-3+\frac{x}{x^2+4}\] if i did my arithmetic right above

OpenStudy (anonymous):

yup yup good arithmetic

OpenStudy (zarkon):

you did :)

myininaya (myininaya):

\[\int\limits_{}^{}(x-3+\frac{x}{x^2+4})dx\]

myininaya (myininaya):

we can use substitution for the 3rd integral \[\int\limits_{}^{}x dx+\int\limits_{}^{}-3 dx+\int\limits_{}^{}\frac{x}{x^2+4} dx\]

myininaya (myininaya):

lol

myininaya (myininaya):

\[\frac{x^2}{2}-3x+\int\limits\limits_{}^{}\frac{x}{x^2+4}dx \]

myininaya (myininaya):

\[u=x^2+4 => du=2x dx =>\frac{1}{2} du=x dx\] \[\frac{x^2}{2}-3x+\frac{1}{2}\int\limits_{}^{} \frac{du}{u}+C=\frac{x^2}{2}-3x+\frac{1}{2}\ln|u|+C\]

myininaya (myininaya):

\[\frac{x^2}{2}-3x+\frac{1}{2}\ln(x^2+4)+C \]

OpenStudy (zarkon):

I like that you dropped the absolute value since you didn't need it in the end :)

myininaya (myininaya):

:) right because x^2+4>0

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (zarkon):

:O

myininaya (myininaya):

i'm puzzled

myininaya (myininaya):

did you not follow anything i did?

OpenStudy (anonymous):

lol

OpenStudy (zarkon):

I would guess no :)

OpenStudy (anonymous):

I need answer now!

OpenStudy (anonymous):

MIA....

OpenStudy (anonymous):

ugh my comp keeps freezing sorry